# Low-Rank Key Value Attention

### v1

@@ 1 Introduction
The design of LRKV differs from previous KV-sharing methods in two respects. First, it retains a full-rank shared component rather than constraining all heads to use the same low-rank representation, which maintains capacity for representing high-dimensional dependencies. Second, the low-rank residuals are trained jointly with the shared projection during pretraining, allowing the model to learn how much variation each head requires. This formulation provides a continuous trade-off between the efficiency of shared KVs and the flexibility of independent projections.
We evaluate LRKV on large-scale language model pretraining and downstream tasks, measuring validation loss, compute-normalized convergence, and KV cache usage across standard MHA, MQA/GQA, and MLA baselines. LRKV consistently improves loss reduction and downstream performance at a given compute budget while reducing KV cache memory by 40-50%. We further provide an operator-space analysis of head diversity showing that LRKV preserves functional head specialization under compression, and characterize the role of residual rank in this trade-off. These findings establish that LRKV’s empirical success stems from aligning with the natural low-rank structure of attention projections while preserving the degrees of freedom necessary for head specialization.
We evaluate LRKV on large-scale language modeling benchmarks and analyze its effect on pretraining dynamics, memory efficiency, and downstream generalization. Across model sizes, LRKV attains comparable perplexity and downstream accuracy to standard MHA while reducing KV cache size by a factor of two to four, depending on configuration. These results suggest that a large portion of per-head variation in KV representations can be captured through low-rank structure, and that explicit factorization offers a principled route to improving the memory efficiency of pretraining.
2 Related Work
@@
Comparison to Multi-Latent Attention. Multi-Latent Attention (MLA) (Liu et al., 2024a) reduces KV cache memory by compressing token representations into a shared low-dimensional latent space before caching. During attention, interaction with the cache therefore requires projecting queries and/or values through this latent bottleneck. This achieves strong memory compression but constrains all heads to operate through the same latent representation, limiting per-head expressivity and restricting positional encoding choices (e.g., requiring partial RoPE).
@@
3 Methodology
Transformers represent each token in a sequence through queries, keys, and values that interact via scaled dot-product attention. For an input matrix $\mathbf{X}\in\mathbb{R}^{L\times d}$, the $h$-th attention head computes
$\mathbf{Q}_{h}=\mathbf{X}\mathbf{W}_{h}^{Q},\qquad\mathbf{K}_{h}=\mathbf{X}\mathbf{W}_{h}^{K},\qquad\mathbf{V}_{h}=\mathbf{X}\mathbf{W}_{h}^{V},$
where $\mathbf{W}_{h}^{Q,K,V}\in\mathbb{R}^{d\times d_{h}}$ and $d_{h}=d/H$ for $H$ heads. The head output is
$\mathbf{O}_{h}=\mathrm{softmax}\!\left(\frac{\mathbf{Q}_{h}\mathbf{K}_{h}^{\top}}{\sqrt{d_{h}}}\right)\mathbf{V}_{h},$
and outputs from all heads are concatenated and projected to form the layer output. During autoregressive decoding, the KV cache stores $\mathbf{K}_{h}$ and $\mathbf{V}_{h}$ for all previous tokens and heads, incurring per-layer memory
$M_{\text{standard}}=2LHd_{h}=2Ld.$
Low-Rank KV Adaptation (LRKV). We parameterize each head’s key/value projection as a shared base plus a head-specific low-rank residual, while keeping the resulting projection shape $d\times d_{h}$:
$\mathbf{W}_{h}^{K}=\mathbf{W}_{\mathrm{shared}}^{K}+\mathbf{U}_{h}^{K}{\mathbf{B}_{h}^{K}}^{\top},\qquad\mathbf{W}_{h}^{V}=\mathbf{W}_{\mathrm{shared}}^{V}+\mathbf{U}_{h}^{V}{\mathbf{B}_{h}^{V}}^{\top},$ (1)
where
$\mathbf{W}_{\mathrm{shared}}^{K,V}\in\mathbb{R}^{d\times d_{h}},\quad\mathbf{U}_{h}^{K,V}\in\mathbb{R}^{d\times r},\quad\mathbf{B}_{h}^{K,V}\in\mathbb{R}^{d_{h}\times r},\quad r\ll d_{h}.$
The effective keys and values are
$\mathbf{K}_{h}=\mathbf{X}\mathbf{W}_{h}^{K},\qquad\mathbf{V}_{h}=\mathbf{X}\mathbf{W}_{h}^{V}.$
When $r=0$, LRKV reduces to complete KV sharing (MQA-style) within a layer; increasing $r$ interpolates toward standard MHA. All parameters in Equation 1 are optimized jointly during pretraining. Gradients from the shared and residual paths are additive, allowing the model to learn how much per-head variation is required. LRKV constrains head-specific projections to an affine family centered at $\mathbf{W}_{\mathrm{shared}}^{K,V}$, providing a continuous spectrum between full sharing ($r=0$) and full independence (large $r$).
LRKV caching scheme. During decoding, the bottleneck is storing $\mathbf{K}_{h},\mathbf{V}_{h}\in\mathbb{R}^{L\times d_{h}}$ for every head. LRKV caches a shared component once per layer and compact per-head latents:
  1. 1.
    Shared features (cached once per layer):
    $\mathbf{K}_{\mathrm{shared}}=\mathbf{X}\mathbf{W}_{\mathrm{shared}}^{K}\in\mathbb{R}^{L\times d_{h}},\qquad\mathbf{V}_{\mathrm{shared}}=\mathbf{X}\mathbf{W}_{\mathrm{shared}}^{V}\in\mathbb{R}^{L\times d_{h}}.$
  2. 2.
    Per-head low-rank latents (cached per head):
    $\mathbf{R}_{h}^{K}=\mathbf{X}\mathbf{U}_{h}^{K}\in\mathbb{R}^{L\times r},\qquad\mathbf{R}_{h}^{V}=\mathbf{X}\mathbf{U}_{h}^{V}\in\mathbb{R}^{L\times r}.$
The full per-head features can be expressed as
$\mathbf{K}_{h}=\mathbf{K}_{\mathrm{shared}}+\mathbf{R}_{h}^{K}{\mathbf{B}_{h}^{K}}^{\top},\qquad\mathbf{V}_{h}=\mathbf{V}_{\mathrm{shared}}+\mathbf{R}_{h}^{V}{\mathbf{B}_{h}^{V}}^{\top}.$ (2)
Attention computation without explicit reconstruction. Naively materializing $\mathbf{K}_{h},\mathbf{V}_{h}$ for all cached tokens would cost $O(Lrd_{h})$. Instead, we exploit associativity to avoid forming $\mathbf{K}_{h},\mathbf{V}_{h}$ explicitly. For a decoding step with query $\mathbf{q}_{h}\in\mathbb{R}^{d_{h}}$,
and for attention weights $\mathbf{a}_{h}\in\mathbb{R}^{1\times L}$,
Equation 3 and Equation 4 enable LRKV to be implemented inside fused attention kernels without reconstructing full per-head KV tensors.
To situate LRKV relative to existing attention mechanisms,  Table 1 summarizes how different methods parameterize key/value projections and where compression is applied.
Method KV projection form Rank constraint Compression axis
$M_{\text{standard}}=2LHd_{h}.$
LRKV stores shared $\mathbf{K}_{\mathrm{shared}},\mathbf{V}_{\mathrm{shared}}$ once per layer plus per-head latents:
$M_{\text{LRKV}}=\underbrace{2Ld_{h}}_{\text{shared }(K,V)}+\underbrace{2LHr}_{\text{per-head latents }(R^{K},R^{V})}=2L(d_{h}+Hr).$ (5)
Thus the memory ratio is
Compute overhead during decoding (FLOPs). Per decoding step, standard attention has dominant per-head computational cost $O(Ld_{h})$ from query-key dot products and value aggregation. Using the distributed forms in Equation 3 and Equation 4, LRKV adds an extra
$\Delta\mathrm{FLOPs}=O(Lr+rd_{h})$
per head per decoding step.
Parameter complexity. Standard per-layer K/V parameters scale as $2Hdd_{h}$. LRKV uses a shared base plus low-rank factors:
which can be lower or higher than standard depending on $(H,r,d_{h})$; in practice we choose $r$ to prioritize KV-cache reduction with minimal quality impact. In our experiments, LRKV remains competitive or better even when $r$ is chosen so that the K/V parameter count is below standard MHA.
4 Shared Subspaces, Spectral Bias, and Information Structure
The MHA mechanism was originally introduced to allow multiple attention “views” of the same sequence, with different heads capturing complementary dependencies (Vaswani et al., 2017). However, empirical analysis of pretrained Transformers reveals that many attention heads are highly correlated and often attend to overlapping tokens or features (Clark et al., 2019; Michel et al., 2019; Bhojanapalli et al., 2021). This redundancy suggests that head-specific projections inhabit a shared, low-dimensional manifold shaped by statistical regularities in the data and by the optimization biases of gradient descent. LRKV can be viewed as an explicit parameterization of this structure, disentangling global and head-specific information within a single architectural formulation.
Shared subspace and spectral structure.

### v2

@@ 4 Experiments
4.1 Pretraining Results
Final pretraining performance.
Table 1 shows LRKV achieves competitive test loss across all scales (128M, 1.2B, 2.5B, 6.3B), with the best performance at 128M and 2.5B scales. At 6.3B scale, LRKV reaches 2.288 CE (0.692 BPB), outperforming MHA (2.319 CE), MQA (2.304 CE), GQA (2.354 CE), and MLA (2.377 CE). At 1.2B scale, LRKV achieves 2.509 CE (0.758 BPB) with only 48% of MHA’s cache—outperforming all baselines while using half the memory. At 2.5B, LRKV achieves 0.719 BPB with only 48.4% of MHA’s cache—a strictly better accuracy-memory tradeoff. Cache efficiency improves at larger scales (52.6% → 48% → 48.4% → 45.1%), making LRKV increasingly attractive for large models.
Table 2: Downstream task performance after midtraining (128M, 2.5B, and 6.3B scales). LRKV achieves the highest combined accuracy across all three scales (18.9%, 37.9%, 40.2%) on five diverse benchmarks, demonstrating that superior pretraining performance translates to stronger downstream capabilities. Combined score (Comb.) is the average across all five benchmarks. HE = HumanEval.
128M 2.5B 6.3B
Model ARC-E ARC-C MMLU GSM8K HE Comb. ARC-E ARC-C MMLU GSM8K HE Comb. ARC-E ARC-C MMLU GSM8K HE Comb.
Standard MHA 26.4 26.0 27.2 1.0 8.8 17.9 66.6 47.1 39.3 10.2 13.4 35.3 72.7 53.6 42.7 10.6 14.6 38.8
GQA 28.0 25.4 25.7 0.5 4.6 16.8 65.4 49.6 40.5 10.0 13.4 35.8 69.0 48.2 40.9 9.9 13.4 36.3
MQA 28.0 27.1 27.8 1.3 2.4 17.3 65.2 47.3 40.2 10.3 3.7 33.3 67.2 47.0 40.5 8.2 4.3 33.4
MLA 27.3 27.8 26.3 1.1 7.7 18.0 67.5 51.5 41.9 10.8 12.8 36.9 69.4 51.3 40.9 9.5 13.4 36.9
LRKV 26.7 30.2 28.4 0.3 9.1 18.9 70.7 53.8 42.2 11.3 11.7 37.9 75.0 58.0 44.5 11.5 12.8 40.2
Cross-scale training efficiency and compute normalization.  Figure 2 presents test performance across all four model scales, measured in cross-entropy loss with dual-axis compute metrics (training tokens and cumulative FLOPs).
LRKV demonstrates competitive performance across scales, achieving the lowest test loss at 128M and 2.5B scales while remaining highly competitive at 1.2B and 6.3B, showing consistent advantages across three orders of magnitude in model size. The top x-axis shows cumulative training compute (ExaFLOPs), accounting for mechanism-specific per-token costs (MQA -6%, GQA -3.2%, MLA -3.7%, LRKV +0.8% at 2.5B scale). [^2] The reported per-token costs reflect inference FLOPs. During training, LRKV materializes full K/V matrices (as do all methods), so training cost per step is comparable to MHA. The 18-25% “training compute savings” refer to sample efficiency (fewer steps to reach target loss), not per-step speedup. When normalized by total compute rather than token count, LRKV maintains its performance advantage: at every ExaFLOP milestone, LRKV achieves lower test loss than all baselines. At 2.5B scale, LRKV reaches equivalent baseline performance 18–30% faster (23.6% average across all methods), while no baseline reaches LRKV’s final performance even after exhausting the full compute budget. This establishes LRKV as dominant in both sample efficiency and final quality. At the largest 6.3B scale, the performance gap widens: LRKV achieves 2.288 CE versus Standard MHA’s 2.319 CE (1.3% improvement), demonstrating that LRKV’s architectural advantages strengthen rather than diminish with scale.
Figure 3: LRKV achieves superior training efficiency alongside best performance (2.5B scale). (a) Memory vs Performance: Test BPB versus KV cache percentage for all methods. LRKV achieves optimal trade-off with lowest BPB at 48.4% cache usage (2.5B scale). (b) Training Efficiency Advantage: LRKV reaches each baseline’s final test loss, quantifying training compute savings. LRKV reaches all baselines’ performance earlier.
Training efficiency analysis. Beyond superior converged performance, LRKV demonstrates remarkable sample efficiency at 2.5B scale (Figure 3). LRKV reaches each baseline’s final validation performance 18-30% faster, averaging 23.6% training compute savings across all baselines while achieving better final performance. Critically, this reveals an asymmetric advantage: LRKV reaches any baseline’s performance target early in training, but no baseline reaches LRKV’s final performance (0.719 BPB) even after the full 50B token budget.
Long Context Pretraining. We extend evaluation to 8192 token sequences using 512M parameter models trained for 50B tokens (153.6 ExaFLOPs). Figure 4 shows that LRKV achieves 2.67 test loss, outperforming MHA (2.74) by 2.7% while using only 48.2% of its KV cache. LRKV also surpasses GQA (2.70, -1.3%), MLA (2.71, -0.9%), and MQA (2.73, -0.4%). Notably, all KV-efficient methods outperform Standard MHA at long context, suggesting compression provides implicit regularization benefits, though LRKV’s architectural advantages remain pronounced. These results validate LRKV’s effectiveness at extended sequence lengths.
Figure 4: Long context pretraining for 512M parameter models. Test cross-entropy loss curves for models trained with 8192-token sequence length on 50B tokens (153.6 ExaFLOPs). LRKV outperforms Standard MHA by 2.7% while using 48.2% cache.
4.2 Downstream Task Performance
To evaluate whether LRKV’s pretraining advantages translate to practical capabilities, we perform supervised midtraining on a diverse instruction-following dataset (568K examples from SmolTalk (Allal et al., 2024), MMLU (Hendrycks et al., 2021), and GSM8K (Cobbe et al., 2021)) and evaluate on five standard benchmarks. Table 2 shows final downstream performance across three scales (128M, 2.5B, 6.3B).
Figure 5: LRKV rank ablation with performance-memory tradeoff analysis (128M, 100B tokens). (a) Final test cross-entropy loss versus KV cache size (relative to Standard MHA) for LRKV rank ablations ($r\in\{8,16,32,64,128\}$), MLA latent dimension ablations ($d\in\{128,192,384\}$), and baselines (Standard MHA, GQA with 3 groups, MQA). LRKV dominates the performance-memory tradeoff space. (b) LRKV training dynamics across ranks, showing consistent convergence and monotonic improvement with increasing rank.
Figure 6: Gauge-invariant head similarity matrices (2.5B scale). Heatmaps show pairwise similarities $s_{ij}=\text{tr}((\mathbf{W}_{i}^{K})^{\top}\mathbf{W}_{j}^{K}(\mathbf{W}_{j}^{Q})^{\top}\mathbf{W}_{i}^{Q})$ normalized by Frobenius norms for all 18 heads. Red indicates high similarity, blue indicates independence. LRKV exhibits similarity structure nearly identical to Standard MHA with predominantly dark off-diagonal regions.
Scale-consistent superiority across task categories. LRKV achieves the highest combined accuracy at all three scales: 18.9% (128M), 37.9% (2.5B), and 40.2% (6.3B), demonstrating consistent advantages across three orders of magnitude in model size. At 2.5B scale, LRKV outperforms Standard MHA (35.3%), GQA (35.8%), MQA (33.3%), and MLA (36.9%), with particularly strong gains on knowledge-intensive and reasoning tasks: ARC-Easy (+4.1pp over MHA), ARC-Challenge (+6.7pp), MMLU (+2.9pp), and GSM8K (+1.1pp). At 6.3B scale, LRKV achieves 40.2% combined versus MHA’s 38.8%, with the largest improvements on reasoning benchmarks (ARC-Challenge: 58.0% vs 53.6%, MMLU: 44.5% vs 42.7%).
Notably, MQA shows catastrophic degradation on code generation (HumanEval: 2.4% at 128M, 3.7% at 2.5B, 4.3% at 6.3B versus 13.4%, 11.7-13.4%, and 12.8-14.6% for other methods), confirming that complete KV sharing particularly harms structured generation tasks requiring precise long-range dependencies. LRKV avoids this pathology through per-head residuals that preserve specialization.
Pretraining quality predicts downstream performance. The strong correlation between pretraining BPB and downstream accuracy (R²=0.786 at 2.5B scale, see Appendix E) confirms that architectural improvements generalizing across pretraining data translate directly to task-specific capabilities. LRKV’s 2.6 percentage point downstream advantage over Standard MHA at 2.5B scale (37.9% vs 35.3%) stems directly from its superior pretraining performance (0.719 vs 0.723 BPB). This validates that LRKV’s low-rank factorization provides fundamental capacity gains that manifest across both language modeling and downstream evaluation, rather than overfitting to pretraining objectives.
4.3 Why Low-Rank Key-Value Attention Works
This subsection provides an empirical analysis explaining why LRKV preserves or exceeds the modeling quality of standard attention despite substantially reducing KV-cache memory. LRKV’s design is motivated by a practical constraint: standard MHA duplicates K/V representations across $H$ heads, causing memory cost to scale linearly with head count. While prior analyses suggest attention heads exhibit substantial redundancy (Michel et al., 2019; Clark et al., 2019), heads also specialize for distinct syntactic and semantic patterns (Zhang et al., 2023), implying that complete KV sharing (MQA) may degrade quality. LRKV addresses this tension through additive factorization: $\mathbf{W}_{h}=\mathbf{W}_{\mathrm{shared}}+\mathbf{U}_{h}\mathbf{B}_{h}^{\top}$, which separates a full-rank shared basis from compact per-head residuals. We now examine three questions: (1) whether appropriate rank selection is critical for quality, (2) whether LRKV preserves architectural head diversity, and (3) whether the learned factorization approaches mathematical optimality.
Figure 7: LRKV preserves head diversity across scales. Gauge-invariant effective rank shows LRKV with sufficient rank matches Standard MHA at 128M, while at 2.5B LRKV achieves 98.3% vs 98.9% for MHA using 48.4% of KV cache.
Analysis setup. For a given Transformer layer, we analyze head-specific projection matrices $\{\mathbf{W}_{h}^{Q},\mathbf{W}_{h}^{K},\mathbf{W}_{h}^{V}\}_{h=1}^{H}$ through their gauge-invariant bilinear forms on the final pretrained checkpoint. For LRKV, we quantify geometric separation using cosine similarity between shared and residual projections, and measure subspace overlap via principal-angle-based metrics. To assess head diversity, we use gauge-invariant similarity metrics based on attention bilinear forms $\mathbf{A}_{h}=\mathbf{W}_{h}^{Q}(\mathbf{W}_{h}^{K})^{\top}$, which are invariant to per-head rotations. We extend this analysis using PCA in bilinear form space by centering the Gram matrix $G$ (where $G_{ij}=\langle A_{i},A_{j}\rangle_{F}$) via the kernel PCA transformation (Smola and Schölkopf, 1998): $G_{\text{centered}}=G-G_{\text{row}}-G_{\text{col}}+G_{\text{mean}}$. This removes the mean bilinear form and reveals the intrinsic dimensionality of head specialization. To assess optimality, we compare LRKV’s learned decomposition to the mathematically optimal rank-$r$ truncated SVD of standard MHA projections, computed post-hoc.
Rank selection determines capacity and performance. We first examine how residual rank $r$ affects modeling quality through a systematic ablation study. Figure 5(b) shows training curves for LRKV with ranks $r\in\{8,16,32,64,128\}$ on 128M parameter models with 100B tokens, demonstrating monotonic improvement as rank increases: $r=128$ achieves the best performance (CE=2.877, BPB=4.156), while $r=8$ shows the worst (CE=2.908, BPB=4.201). The performance gap of approximately 1.06% confirms that residual rank is a critical capacity control. Figure 5(a) contextualizes this ablation by plotting final performance against KV cache size for all evaluated methods, revealing that LRKV achieves superior performance-memory tradeoffs across the rank spectrum compared to MLA latent dimension ablations and baselines.
The constrained $r=16$ configuration underperforms standard MHA (0.881 vs 0.878 BPB at 128M), lacking capacity to capture head-specific variation. In contrast, sufficient rank ($r\geq 64$) outperforms MHA (0.875 BPB at 128M, 0.719 at 2.5B vs 0.878 and 0.723), demonstrating that representational capacity is the limiting factor. In our experiments, we find that $r\approx 0.36$-$0.43\times d_{h}$ provides the necessary capacity for LRKV to exceed standard attention performance and MQA, GQA and MLA baselines (see Appendix G for scale-specific rank values). This empirical regularity across scales suggests that this rank range aligns with the intrinsic rank of attention projections while optimizing the performance-memory tradeoff (see  Appendix B for spectral analysis).
LRKV preserves functional head diversity. Figure 6 confirms LRKV exhibits nearly identical similarity patterns to Standard MHA. Quantitatively, we measure head diversity using gauge-invariant metrics based on attention bilinear forms $A_{h}=W_{h}^{Q}(W_{h}^{K})^{\top}$, computing effective rank via eigenvalue entropy ( Appendix B). LRKV (r=64) achieves 98.3% effective rank at 2.5B scale versus 98.9% for Standard MHA—a negligible 0.6pp difference (Figure 7). In contrast, MQA achieves only 86.2% and GQA 95.4%.
Interpreting uncentered vs. PCA-based effective rank. The distinction between uncentered (98.3%) and PCA-based (93.5%) effective rank reveals LRKV’s factorization structure. Uncentered analysis measures total variance including the shared mean direction—the global structure captured by $\mathbf{W}_{\text{shared}}$. PCA-based analysis centers the Gram matrix, isolating variance around this mean and measuring true head independence. The modest 4.8pp gap indicates LRKV achieves diversity primarily through genuine per-head specialization rather than merely perturbing a dominant shared structure. For comparison, MQA shows dramatic improvement from uncentered (86.2%) to centered (91.0%)—a compensation effect where forced KV sharing creates a strong mean direction, but heads recover diversity by aggressively diversifying query projections around this baseline (see subsection C.4 for detailed analysis).
Figure 8: Magnitude scaling in LRKV is absorbed by post-projection normalization. Frobenius norms of key projections comparing standard MHA with LRKV’s shared, residual, and total (shared + residual) components for (a) 128M and (b) 2.5B models. LRKV projections operate in a higher-magnitude regime than standard MHA, with both shared and residual components individually exceeding standard MHA magnitudes.
PCA-based analysis reveals compensation mechanisms. Applying PCA in bilinear form space ( subsection C.4), LRKV achieves 93.5% PCA-based effective rank at 2.5B versus 94.0% for Standard MHA—within 0.5% despite rank-64 factorization. Remarkably, PCA reveals a compensation effect in MQA: its centered effective rank (91.0%) increases versus uncentered (86.2%), opposite to other architectures. This indicates MQA heads compensate for forced KV sharing by diversifying query projections more aggressively.
[…]
Why magnitude scaling doesn’t degrade quality. The higher-magnitude regime emerges from unconstrained optimization of the additive factorization: gradients can independently scale $\mathbf{W}_{\text{shared}}$ and residuals $\mathbf{U}_{h}\mathbf{B}_{h}^{\top}$ without affecting their sum’s direction. RMSNorm applied after projection normalizes representations before attention computation, making the attention mechanism operate on directional information (cosine similarity) rather than absolute magnitudes. This architectural property ensures that magnitude scaling affects only the internal parameterization, not the functional behavior. The magnitude difference serves as a diagnostic: it indicates that optimization successfully allocated capacity between shared and residual pathways, with both contributing substantively to the final projection rather than one dominating.
5 Discussion
The empirical analyses collectively reveal why LRKV achieves superior performance despite KV cache reduction:
(1) Appropriate rank selection is critical. The range $r\approx 0.36$-$0.43\times d_{h}$ (46-55 for $d_{h}=128$) provides sufficient degrees of freedom for per-head specialization while constraining the factorization to exploit structured redundancy. Below this threshold (e.g., $r=16$, 0.881 BPB), capacity becomes limiting; above it ($r\geq 64$, 0.875 BPB), performance exceeds MHA, confirming that representational capacity (not geometric properties) determines quality.
(2) LRKV preserves architectural head diversity. PCA-based analysis shows LRKV achieves 93.5% effective rank versus 94.0% for standard MHA at 2.5B scale (Figure 7)—within 0.5% despite rank-64 factorization and 48.4% cache size. This near-perfect preservation validates that low-rank residuals provide sufficient capacity for heads to occupy independent dimensions in bilinear form space. The consistency across all 36 layers demonstrates depth-invariant factorization quality.
(3) Compensation mechanisms explain baseline behavior. The PCA-based methodology reveals phenomena invisible to prior metrics: MQA’s 4.8pp improvement from uncentered to centered effective rank quantifies query compensation - heads recover diversity by specializing queries around forced shared KVs. This explains why MQA remains viable despite complete KV sharing, and positions LRKV’s approach (preserving KV and query diversity) as superior.
[…]
These properties collectively enable LRKV to achieve strong pretraining performance (0.692 BPB at 6.3B) and downstream accuracy (40.2% combined at 6.3B) while reducing KV cache to 45-53% of standard attention. The analyses confirm LRKV exploits structured redundancy without sacrificing the head specialization that makes MHA effective.
6 Conclusion
We introduced Low-Rank Key-Value attention, which decomposes key/value projections into a shared dense component and compact per-head low-rank residuals. LRKV achieves (1) 45-53% KV cache reduction while attaining the best pretraining loss (0.692 BPB at 6.3B) and 18-25% training compute savings; (2) the highest downstream performance across 5 benchmarks for varying model sizes; and (3) near-complete preservation of head diversity give sufficiently lower rank (93.5% vs. 94% for standard MHA), confirming that LRKV exploits structured redundancy without sacrificing specialization. Our gauge-invariant analysis and rank ablations show that residual rank $r\approx 0.36$-$0.43\times d_{h}$ provides a critical threshold for preserving head specialization while enabling substantial KV cache reduction, positioning LRKV as a practical drop-in replacement for standard attention under memory constraints.
Appendix A Related Work (Full Version)
@@
Comparison to Multi-Latent Attention. Multi-Latent Attention (MLA) (Liu et al., 2024) reduces KV cache memory by compressing token representations into a shared low-dimensional latent space before caching. During attention, interaction with the cache therefore requires projecting queries and/or values through this latent bottleneck. This achieves strong memory compression but constrains all heads to operate through the same latent representation, limiting per-head expressivity and restricting positional encoding choices (e.g., requiring partial RoPE).
@@
Appendix B Shared Subspaces, Spectral Bias, and Information Structure